# On the optimality of coin-betting for mean estimation

### v1

@@ 4 Optimality of the coin-betting e-class
Now, to check that $F$ majorises all the e-variables, fix $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$ and $x\in[0,1]$. Consider a $P_{x}\in\mathcal{H}_{\mu}$ with mass $1/F_{\mu}(x)$ on $x$. Then, $1\geq\langle P_{x},E\rangle\geq P_{x}(x)E(x)=E(x)/F_{\mu}(x)$, and we conclude. $\qed$
Theorem 1
@@
First, let us show that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class. Fix an arbitrary $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$. Define the sets
$B_{0}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in[0,\mu]}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\qquad\text{and}\qquad B_{1}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in[\mu,1]}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\,,$
@@
where we recall that $I_{\mu}=[(1-\mu)^{-1},\mu^{-1}]$ and $E_{\beta}:x\mapsto 1+\beta(x-\mu)$. Both sets are closed and convex (as they are intersections of closed and convex sets). By Lemma 5, $E_{(\mu-1)^{-1}}\in B_{0}$ and $E_{\mu^{-1}}\in B_{1}$, so $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu]$, for some $\beta_{0}$ and $\beta_{1}$ in $I_{\mu}$. We will now show that $B_{0}\cap B_{1}\neq\varnothing$, or equivalently that $\beta_{0}\geq\beta_{1}$. Assume that this was not the case and $\beta_{0}<\beta_{1}$. Let $\beta^{\star}\in(\beta_{0},\beta_{1})$. Then, $\beta^{\star}\notin B_{0}$ and $\beta^{\star}\notin B_{1}$. In particular, there are $x_{0}\in[0,\mu]$ and $x_{1}\in[\mu,1]$ such that $E(x_{0})>E_{\beta^{\star}}(x_{0})$ and $E(x_{1})>E_{\beta^{\star}}(x_{1})$. Notice that since $\delta_{\mu}\in\mathcal{H}_{\mu}$, it must be that $E(\mu)\leq 1$, and so $x_{0}<\mu$ and $x_{1}>\mu$. Hence, $\mu\in(x_{0},x_{1})$ and so there is $\hat{P}\in\mathcal{H}_{\mu}$ with support $\{x_{0},x_{1}\}$. Notice that $\langle\hat{P},E_{\beta^{\star}}\rangle=1+\beta^{\star}(\langle\hat{P},X\rangle-\mu)=1$. But $E$ is strictly larger than $E_{\beta^{\star}}$ on $\mathrm{Supp}(\hat{P})$, and so $\langle\hat{P},E\rangle>1$, which is a contradiction since $E$ is an e-variable. We conclude that $\beta_{0}\geq\beta_{1}$ and so there exists $\hat{\beta}\in B_{0}\cap B_{1}$. By construction, $E_{\hat{\beta}}\in\mathcal{E}^{\mathrm{cb}}_{\mu}$ and $E_{\hat{\beta}}\succeq E$. Because the choice of $E$ was arbitrary, $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class.
@@
Now that we have established that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class, by Lemma 4 we only need to show that all its elements are maximal to conclude that it is optimal. Fix $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, and consider an e-variable $\hat{E}\in\mathcal{E}_{\mathcal{H}_{\mu}}$ such that $\hat{E}\succeq E$. Fix any $x\in[0,1]$. By Lemma 5, there is $P\in\mathcal{H}_{\mu}$ such that $P(\{x\})>0$. Since $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, we have $\langle P,E\rangle=1$, and so $0\leq P(\{x\})(\hat{E}(x)-E(x))\leq\langle P,\hat{E}-E\rangle=\langle P,\hat{E}\rangle-1\leq 0$. Since $P(\{x\})>0$, we get $\hat{E}(x)=E(x)$ and, $x$ being arbitrary, $\hat{E}=E$. Hence $E$ is maximal, as it has no strict majoriser. $\qed$
@@ 5 Perspectives
Theorem 1 gives a rigorous sense to the claim that the coin-betting formulation is optimal among the e-variable-based approaches to build confidence sequences for $\mu^{\star}$. To formalise this assertion, we introduced the notions of majorising and optimal e-classes, which, in the author’s view, are of independent interest in the context of sequential testing, beyond the scope of this letter. The main novelty of these concepts lies in defining “optimality” in terms of e-classes, rather than individual e-variables. This perspective contrasts with the notion of log-optimality of a single e-variable with respect to an alternative hypothesis, widely discussed in the literature (e.g., Koolen and Grünwald, 2022; Grünwald et al., 2024; Larsson et al., 2024).
Rather than log-optimality, what probably best matches our way of “ranking” e-variables is the idea of admissibility introduced by Ramdas et al. (2022). Roughly speaking, let us say that a wealth-process (namely, a non-negative adapted process starting at one that is a super-martingale under every $P\in\mathcal{H}$) is admissible if it is not dominated by any other wealth-process. This concept is tightly related to our notion of maximal e-variables, as it is not hard to see that any wealth-process $(M_{n})_{n\geq 1}$ on $\mathcal{H}_{\mu}$ is admissible if, and only if, it can be associated with a sequence $(E_{t})_{t\geq 1}$ of maximal e-variables such that $M_{n}=\prod_{t=1}^{n}E_{t}(X_{t}^{\prime})$, with $(X_{t}^{\prime})_{t\geq 1}$ a sequence generated by some $P\in\mathcal{H}_{\mu}$. Thus, in our setting, Theorem 1 provides a characterisation of the set of all admissible wealth-processes for $\mathcal{H}_{\mu}$. Let us emphasise that the tools from Ramdas et al. (2022) alone are insufficient to achieve this, as the necessary conditions for admissibility provided by their work rely on the existence of a “reference measure” for the hypothesis, [^3] A hypothesis $\mathcal{H}$ admits a “reference measure” if there exists a Borel measure $Q$ such that $P\ll Q$ for all $P\in\mathcal{H}$. a property that $\mathcal{H}_{\mu}$ does not satisfy. More precisely, their results are enough to prove that every e-variable in the coin-betting e-class is maximal, but not that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class. This point is subtle but essential: establishing the optimality of coin-betting requires studying the properties of a set of e-variables (or equivalently, wealth-processes) rather than focusing on single e-variables or wealth-processes individually. Crucially, it is precisely the fact that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class that ensures that nothing is lost by relying on the coin-betting approach.
@@
To the author’s knowledge, this is the first work to rigorously examine the optimality of the coin-betting formulation, as previous discussion on the topic has relied solely on heuristic arguments. For instance, Waudby-Smith and Ramdas (2023) justify restricting to the coin-betting e-class by noticing that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is precisely the set ¯$\bar{\mathcal{E}}_{\mathcal{H}_{\mu}}$ of e-variables whose expectation equals one under every $P\in\mathcal{H}_{\mu}$. However, this property does not generally imply the optimality of the e-class, as it merely indicates that all its elements are maximal. As a simple counterexample, consider $\mathcal{H}^{\prime}_{\mu}=\{P\in\mathcal{P}\,:\,\langle P,X\rangle\leq\mu\}$. Then, ¯$\bar{\mathcal{E}}_{\mathcal{H}^{\prime}_{\mu}}=\{1\}$, while the optimal e-class exists and consists of all functions in the form $E_{\beta}:x\mapsto 1+\beta(x-\mu)$, with $\beta\in[0,\mu^{-1}]$.
To simplify the exposition, we have focused on random variables valued in $[0,1]$. However, all the results presented can be extended to any bounded closed interval. More broadly, the author anticipates that similar conclusions hold for more general settings, such as estimating the mean of bounded vector-valued random variables. Addressing this case requires more advanced analytic tools and lies beyond the scope of this letter.
@@
As a final remark, this paper does not aim to delve into the design of effective coin-betting strategies for obtaining tight confidence sequences. For this, we refer the interested reader to the thorough analysis and discussion by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023).

### v2

@@ 4 Optimality of the coin-betting e-class
[…]
Figure 1: Pictorial representation of the main step in the proof of Theorem 1. $1+\beta_{1}(x-\mu)$ dominates $E$ for $x\in[0,\mu)$, while $1+\beta_{0}(x-\mu)$ dominates $E$ for $x\in(\mu,1]$. If there is $\beta^{\star}\in(\beta_{0},\beta_{1})$, then we can find $u_{0}\in[0,\mu)$ and $u_{1}\in(\mu,1]$ such that $E(u_{0})>1+\beta^{\star}(u_{0}-\mu)$ and $E(u_{1})>1+\beta^{\star}(u_{1}-\mu)$, represented by the points $U_{0}$ and $U_{1}$ being above the purple line $1+\beta^{\star}(x-\mu)$. The probability measure $\hat{P}$ supported on $\{u_{0},u_{1}\}$ with mean $\mu$ is in $\mathcal{H}_{\mu}$. We have $\mathbb{E}_{\hat{P}}[E]>1$. Indeed, this expected value corresponds to the vertical coordinate of the point $U^{\star}$, the intersection of the line connecting $U_{0}$ and $U_{1}$ with the vertical line at $x=\mu$. This is a contradiction if $E$ is an e-variable, in which case it must be that $\beta_{0}\geq\beta_{1}$.
@@
Now, to check that $F_{\mu}$ majorises all the e-variables, fix $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$ and $x\in[0,1]$. Consider a $P_{x}\in\mathcal{H}_{\mu}$ with mass $1/F_{\mu}(x)$ on $x$. Then, $1\geq\mathbb{E}_{P_{x}}[E]\geq P_{x}(x)E(x)=E(x)/F_{\mu}(x)$, and we conclude. ∎
Theorem 1.
@@
First, let us show that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class. For a pictorial representation of this part of the proof, the reader is invited to look at Figure 1. Fix an arbitrary $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$. Define the sets
$B_{0}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in[0,\mu)}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\quad\text{and}\quad B_{1}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in(\mu,1]}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\,,$
@@
where we recall that $I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$ and $E_{\beta}:x\mapsto 1+\beta(x-\mu)$. Both sets are closed and convex (as they are intersections of closed and convex sets). By Lemma 4, $(\mu-1)^{-1}\in B_{0}$ and $\mu^{-1}\in B_{1}$, so $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu^{-1}]$, for some $\beta_{0}$ and $\beta_{1}$ in $I_{\mu}$. We will now show that $B_{0}\cap B_{1}\neq\varnothing$, or equivalently that $\beta_{0}\geq\beta_{1}$. Assume that this was not the case and $\beta_{0}<\beta_{1}$. Let $\beta^{\star}\in(\beta_{0},\beta_{1})$. Then, $\beta^{\star}\notin B_{0}$ and $\beta^{\star}\notin B_{1}$. In particular, there are $u_{0}\in[0,\mu)$ and $u_{1}\in(\mu,1]$ such that $E(u_{0})>E_{\beta^{\star}}(u_{0})$ and $E(u_{1})>E_{\beta^{\star}}(u_{1})$. As $\mu\in(u_{0},u_{1})$, there is $\hat{P}\in\mathcal{H}_{\mu}$ with support $\{u_{0},u_{1}\}$. Note that $\mathbb{E}_{\hat{P}}[E_{\beta^{\star}}]=1+\beta^{\star}(\mathbb{E}_{\hat{P}}[X]-\mu)=1$. But $E$ is strictly larger than $E_{\beta^{\star}}$ on $\mathrm{Supp}(\hat{P})$, and so $\mathbb{E}_{\hat{P}}[E]>1$, which is a contradiction since $E$ is an e-variable. We conclude that $\beta_{0}\geq\beta_{1}$ and so there exists $\hat{\beta}\in B_{0}\cap B_{1}$. By construction, $E_{\hat{\beta}}\in\mathcal{E}^{\mathrm{cb}}_{\mu}$ and $E_{\hat{\beta}}(x)\geq E(x)$ for all $x\in[0,1]$ different from $\mu$. For $x=\mu$, by Lemma 4 we have $E(\mu)\leq F_{\mu}(\mu)=1=E_{\hat{\beta}}(\mu)$, and so $E\preceq E_{\hat{\beta}}$. As the choice of $E$ was arbitrary, $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class.
@@
Now that we have established that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class, by Lemma 3 we only need to show that all its elements are maximal to conclude that it is optimal. Fix $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, and consider an e-variable $\hat{E}\in\mathcal{E}_{\mathcal{H}_{\mu}}$ such that $\hat{E}\succeq E$. Fix any $x\in[0,1]$. By Lemma 4, there is $P\in\mathcal{H}_{\mu}$ such that $P(\{x\})>0$. Since $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, we have $\langle P,E\rangle=1$, and so $0\leq P(\{x\})(\hat{E}(x)-E(x))\leq\mathbb{E}_{P}[\hat{E}-E]=\mathbb{E}_{P}[\hat{E}]-1\leq 0$. Since $P(\{x\})>0$, we get $\hat{E}(x)=E(x)$ and, $x$ being arbitrary, $\hat{E}=E$. Hence $E$ is maximal, as it has no strict majoriser. ∎
5 Optimality beyond single-round e-variables
Up to this point, we have focused on hypotheses defined as subsets of the space $\mathcal{P}_{[0,1]}$, which cannot capture dependencies across multiple rounds. In such setting, the dependency structure among observations is fixed and assumed a priori, rather than being subject to testing. Yet, e-variable-based testing naturally extends to more general hypotheses that span multiple rounds and can account for sequential or dependent data structures. In the second part of the paper, we turn our attention to this richer framework. Remarkably, a slight adaptation of the proof strategy used for Theorem 1 and depicted in Figure 1 allows us to show that coin-betting-based testing is also optimal in the setting where the sequence of observations has a fixed conditional mean $\mu$. We will first show that this is the case with multi-round e-variables for a fixed time horizon, and then extend the result to testing with e-processes. Conversely, we will show that when testing the i.i.d. assumption, the coin-betting approach no longer yields the optimal e-class.
For technical reasons (primarily related to measurability, see discussion in Section 7), for the rest of this work we restrict ourselves to the case of a random variable taking values in a finite (known) set. Therefore, henceforth we let $\mathcal{X}\subseteq\mathbb{R}$ denote a finite set. For simplicity, we assume that $\min\mathcal{X}=0$ and $\max\mathcal{X}=1$, although all our results extend readily to any case where the minimum and maximum values are finite. We note that since $\mathcal{X}$ has finite cardinality, every function $\mathcal{X}^{t}\to\mathbb{R}$ is automatically Borel measurable.
5.1 Optimal e-class for multi-round e-variables
For any $\mu\in(0,1)$ and $T\geq 1$, we let
$\mathcal{H}_{\mu}^{T}=\big{\{}P\in\mathcal{P}_{\mathcal{X}^{T}}\,:\,\mathbb{E}_{P}[X_{t}|X^{t-1}]=\mu\,,\;\forall t\in[1:T]\big{\}}\,,$
where $\mathbb{E}_{P}[X_{1}|X^{0}]=\mathbb{E}_{P}[X_{1}]$. Note that $\mathcal{H}_{\mu}^{T}$ is a hypothesis on $\mathcal{X}^{T}$. As such, its set of e-variables will consist of Borel functions from $\mathcal{X}^{T}$ to $\mathbb{R}$. We denote the set of all e-values relative to $\mathcal{H}_{\mu}^{T}$ as $\mathcal{E}_{\mu}^{T}$.
In a coin-betting game, the wealth of the player at round $T$ is a non-negative function of the observations, which takes the form $M_{T}(x^{T})=\prod_{t=1}^{n}E_{t}(x_{t})$, where $E_{t}$ is in the form $E_{\lambda_{t}}$, with $\lambda_{t}$ chosen as a function of the past observations $x^{t-1}$. It is straightforward to check that under the hypothesis $\mathcal{H}_{\mu}^{T}$, the expectation of $M_{T}$ is always exactly one ($M_{T}$ is a martingale), which shows that $M_{T}$ is an e-variable. Next, we show that the e-class of e-variables in this form coincides is the optimal e-class for $\mathcal{H}_{\mu}^{T}$.
We let $\Lambda^{T}_{\mu}$ be a set of $T$-tuples of functions defined as
$\Lambda^{T}_{\mu}=\big{\{}\lambda^{T}=(\lambda_{1},\dots,\lambda_{T})\,:\,\text{$\lambda_{t}$ is a function from $\mathcal{X}^{t-1}$ to $I_{\mu}$}\big{\}}\,.$
Given $\lambda^{T}\in\Lambda^{T}_{\mu}$, we define the e-variable $E_{\lambda^{T}}\in\mathcal{E}^{T}_{\mu}$ as
$E_{\lambda^{T}}(x^{T})=\prod_{t=1}^{T}\big{(}1+\lambda_{t}(x^{t-1})(x_{t}-\mu)\big{)}\,.$
We define the T round coin-betting e-class $\mathcal{E}^{\mathrm{cb},T}_{\mu}=\{E_{\lambda^{T}}\,:\,\lambda^{T}\in\Lambda_{\mu}^{T}\}$.
Theorem 2.
Fix any $T\geq 1$ and $\mu\in(0,1)$. $\mathcal{E}^{\mathrm{cb},T}_{\mu}$ is the optimal e-class for $\mathcal{H}_{\mu}^{T}$.
The proof of Theorem 2 shares many similarity with that of Theorem 1. In particular, we use the same technique to establish the maximality of e-variables of the form $E_{\lambda^{T}}$. A geometric argument analogous to the one illustrated in Figure 1 is then employed, as part of an induction recursion, to prove that the e-class of interest is majorising. However, the detailed proof being somewhat lengthy, we defer it to the appendix.
5.2 A remark on the i.i.d. case
From what we have shown so far, in a $T$ round game coin-betting yields all and only the maximal e-variables, when testing the hypothesis that the conditional mean is some fixed $\mu\in(0,1)$. Yet, this is not any more the case if we consider a more restrictive hypotheses, stating that the observations are i.i.d. and with mean $\mu$. More concretely, let
^$\hat{\mathcal{H}}_{\mu}^{T}=\big{\{}P=Q^{\otimes T}\,:\,Q\in\mathcal{H}_{\mu}\big{\}}\,$
[…]
To give a concrete example of the mismatch between the e-variables for the conditional and independent scenarios, we consider the “highly symmetrical” case of $\mathcal{X}=\{0,1/2,1\}$, with $\mu=1/2$ and $T=2$. We show at the end appendix that the e-variables in this setting are the functions $E:\{0,1/2,0\}^{2}\to[0,+\infty)$ that satisfy
$\xi_{0}\leq 1\,;\qquad\xi_{1}\leq 1+\sqrt{(1-\xi_{0})(1-\xi_{2})}\,;\qquad\xi_{2}\leq 1\,,$
where $\xi_{0}=E(1/2,1/2)$, $\xi_{1}=(E(1,1/2)+E(1/2,1)+E(1/2,0)+E(0,1/2))/4$, and $\xi_{2}=(E(1,1)+E(1,0)+E(0,1)+E(0,0))/4$. Clearly, the optimal e-class here consists of those e-variables satisfying the constraint on $\xi_{1}$ with equality. We note that all the e-variables in $\mathcal{E}_{1/2}^{T}$ satisfy $\xi_{0}=\xi_{1}=\xi_{2}=1$. As expected, these are indeed maximal e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$, but there are maximal e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$ that are not in $\mathcal{E}_{1/2}^{T}$. The function that equals $4$ on $x^{2}=(1,1/2)$, and 00 everywhere else, is an example of a (maximal) e-variable for ^$\hat{\mathcal{H}}_{1/2}^{2}$, which is not an e-variable for $\mathcal{H}_{1/2}^{2}$.
We leave as a non-trivial open question to characterise the e-variables for ^$\hat{\mathcal{H}}^{T}_{\mu}$ in the general case.
6 E-processes for testing the fixed conditional mean
A key advantage of many testing procedures involving e-variables is anytime validity. One way to achieve this is through the round-wise testing-by-betting approach, which gives rise to a super-martingale (cf. Lemma 1). However, as we have already pointed out, this is not the most general approach. Rather than only focusing on super-martingales, one can instead use any non-negative process whose expectation is upper bounded by one, regardless the stopping time. These are known as e-processes, a generalization of super-martingales, which can enable powerful tests even in scenarios where martingale-based methods lack power (Ramdas et al., 2022b).
As in the previous section, here $\mathcal{X}$ denotes a finite subset of $[0,1]$, with $\min\mathcal{X}=0$ and $\max\mathcal{X}=1$. We denote as $\mathcal{X}^{\infty}$ the space of sequences $(x_{t})_{t\geq 1}\subseteq\mathcal{X}$. For convenience, we will sometimes write $x^{\infty}$ to mean $(x_{t})_{t\geq 1}$. We endow the set $\mathcal{X}^{\infty}$ with the product sigma-field generated by cylinder sets, and we denote as $\mathcal{P}_{\mathcal{X}^{\infty}}$ the space of probability measures on $\mathcal{X}^{\infty}$. For a fixed $\mu\in(0,1)$, in this section we will focus on the hypothesis class
$\mathcal{H}_{\mu}^{\infty}=\big{\{}P\in\mathcal{P}_{\mathcal{X}^{\infty}}\,:\,\mathbb{E}_{P}[X_{t}|X^{t-1}]=\mu\,,\;\forall t\geq 1\big{\}}\,.$
This time, we will not aim to study the e-variables for $\mathcal{H}^{\infty}_{\mu}$, as these would be functions that take a whole sequence as argument. Conversely, we are interested in tools that allow us to stop the test if we think enough data-points have been observed to decide whether or not the hypothesis has to be rejected. To make this rigorous we first need some definitions. First, we call a finite stopping time a function $\tau:\mathcal{X}^{\infty}\to\mathbb{N}$ such that, whenever for some $x^{\infty}$ and $y^{\infty}$ in $\mathcal{X}^{\infty}$ and some $t\in\mathbb{N}$ we have $\tau(x^{\infty})=\tau(y^{\infty})=t$, then $x^{t}=y^{t}$. [^8] It is straightforward to verify that in our setting this definition is equivalent to the perhaps more classical notion of finite e-process as a $\tau$ such that the sets $\{\tau\leq t\}$ are $\mathcal{F}_{t}$ measurable. We say that $\tau$ is bounded by $T\geq 0$ if its image is contained in $[0:T]$. We let $\mathcal{T}_{T}$ denote the set of all stopping times bounded by $T$.
Definition 6.
An e-process (for $\mathcal{H}_{\mu}^{\infty}$) is a sequence $(E_{t})_{t\geq 0}$, of non-negative functions $E_{t}:\mathcal{X}^{t}\to[0,+\infty)$, such that, for every $P\in\mathcal{H}^{\infty}_{\mu}$, any $T\geq 0$ and any $\tau\in\mathcal{T}_{T}$,
$\mathbb{E}_{P}[E_{\tau}]\leq 1\,.$
We denote as $\mathcal{E}_{\mu}^{\infty}$ the set of all the e-processes for $\mathcal{H}_{\mu}^{\infty}$. We call e-process class any subset of $\mathcal{E}_{\mu}^{\infty}$. Similarly to what we have done for the e-variables, we say that the e-process $E$ majorises the e-process $E^{\prime}$ (we write $E\succeq E^{\prime}$) if for every $t\geq 0$ we have $E_{t}\succeq E^{\prime}_{t}$. We say that an e-process class $\mathcal{E}$ is majorising if for every $E\in\mathcal{E}^{\infty}_{\mu}$, there is $E^{\prime}\in\mathcal{E}$ such that $E\preceq E^{\prime}$. We say that $\mathcal{E}$ is optimal if it is majorising and is contained in every other majorising e-process class. We call an e-process $E$ maximal if it is not strictly majorised by any other e-process. With the same reasoning as in Lemma 3, one can show that the optimal e-process class exists if, and only if, the class of all the maximal e-processes is majorising. In such case the optimal class is unique and coincides precisely with all and only the majorising e-processes.
We let $\Lambda_{\mu}^{\infty}$ be the set of all sequences $\lambda^{\infty}=(\lambda_{t})_{t\geq 1}$, such that $\lambda_{t}:\mathcal{X}^{t-1}\to I_{\mu}$. For any $\lambda^{\infty}\in\Lambda_{\mu}^{\infty}$, we define $E^{\lambda^{\infty}}=(E^{\lambda^{\infty}}_{T})_{T\geq 0}$ via
$E^{\lambda^{\infty}}_{T}(x^{T})=E_{\lambda^{T}}(x^{T})=\prod_{t=1}^{T}\big{(}1+\lambda_{t}(x^{t-1})(x_{t}-\mu)\big{)}\,,$
for $T\geq 1$, and $E_{0}=1$. These are the e-processes associated to the wealth of a coin-betting player. As $E^{\lambda^{\infty}}$ defines a martingale under $\mathcal{H}_{\mu}^{\infty}$, it is also an e-process for such hypothesis. We define the coin-betting e-process class $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}=\{E^{\lambda^{\infty}}\,:\,\lambda^{\infty}\in\Lambda^{\infty}_{\mu}\}$.
The next result shows that for $\mathcal{H}_{\mu}^{\infty}$ the coin betting approach is optimal also at the level of e-processes. The proof (reported in the appendix) always follows the ideas introduced in the proof of Theorem 1.
Theorem 3.
For any $\mu\in(0,1)$, $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is the optimal e-process class for $\mathcal{H}_{\mu}^{\infty}$.
@@ 7 Perspectives
Theorem 1 gives a rigorous sense to the claim that the coin-betting formulation is optimal among the e-variable-based approaches to build confidence sequences for $\mu^{\star}$. Theorem 2 and Theorem 3 extended this optimality result to testing the fixed conditional mean of a sequence of bounded observations. To formalise these claims, we introduced the notions of majorising and optimal e-classes, which, in the author’s view, are of independent interest in the context of sequential testing, beyond the scope of this paper. The main novelty of these concepts lies in defining “optimality” in terms of e-classes, rather than individual e-variables. This perspective contrasts with the notion of log-optimality of a single e-variable with respect to an alternative hypothesis, widely discussed in the literature (e.g., Koolen and Grünwald, 2022; Grünwald et al., 2024; Larsson et al., 2024).
The concept of maximality for e-variables, as introduced in this paper, is essentially equivalent to the classical statistical notion of admissibility. An e-variable is considered maximal if no other e-variable strictly dominates it. Likewise, the idea of a majorising e-class corresponds to the classical concept of a complete class, with the optimal e-class representing the minimal complete class of e-variables. For a detailed discussion of these ideas in traditional statistics we refer to Lehmann and Romano (2022). In this work, however, we adopt terminology from partially ordered set (poset) theory to highlight the fact that with e-variables these properties follows from the classical partial order of dominance among functions.
The concept of admissibility in the context of e-variables and e-processes was previously introduced by Ramdas et al. (2022a), whose definition is closely aligned with our notion of maximality. The key distinction lies in the type of dominance considered: their framework often relies on almost sure dominance, whereas our definition of maximality requires everywhere dominance. [^9] The two definitions becomes of course equivalent when $\mathcal{X}$ is finite, but they are not in the case of $\mathcal{X}=[0,1]$, which is the setting of our analysis on single-round e-variables. This stricter requirement is motivated by the will of developing a theory that remains valid even in the absence of a known alternative, where any value in $\mathcal{X}$ might occur at the next observation. Although Ramdas et al. (2022a) provide some necessary and sufficient conditions for admissibility, we emphasize that their results are insufficient to establish the optimality of the coin-betting e-class. Specifically, their approach assumes the existence of a reference measure for the hypothesis, [^10] A hypothesis $\mathcal{H}$ admits a ”reference measure” if there exists a Borel measure $Q$ such that $P\ll Q$ for all $P\in\mathcal{H}$. a condition not satisfied by $\mathcal{H}_{\mu}$. More precisely, their results can show that each individual e-variable in the coin-betting e-class is maximal, but not that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ forms a majorising e-class. This distinction, though subtle, is crucial: proving the optimality of coin-betting requires analyzing the collective properties of a set of e-variables (or equivalently, wealth processes), rather than evaluating them in isolation. It is precisely the fact that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class that guarantees nothing is lost by relying on the coin-betting approach. This point is subtle but essential: establishing the optimality of coin-betting requires studying the properties of a set of e-variables (or equivalently, wealth-processes) rather than focusing on single e-variables or wealth-processes individually. Crucially, it is precisely the fact that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class that ensures that nothing is lost by relying on the coin-betting approach.
@@
To the author’s knowledge, this is the first work to rigorously examine the optimality of the coin-betting formulation, as previous discussion on the topic has relied solely on heuristic arguments. For instance, Waudby-Smith and Ramdas (2023) justify restricting to the coin-betting e-class by noticing that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is precisely the set ¯$\bar{\mathcal{E}}_{\mathcal{H}_{\mu}}$ of e-variables whose expectation equals one under every $P\in\mathcal{H}_{\mu}$. However, this property does not generally imply the optimality of the e-class, as it merely indicates that all its elements are maximal. As a simple counterexample, consider $\mathcal{H}^{\prime}_{\mu}=\{P\in\mathcal{P}\,:\,\langle P,X\rangle\leq\mu\}$. Then, ¯$\bar{\mathcal{E}}_{\mathcal{H}^{\prime}_{\mu}}=\{1\}$, while the optimal e-class exists and consists of all functions in the form $E_{\beta}:x\mapsto 1+\beta(x-\mu)$, with $\beta\in[0,\mu^{-1}]$. Although one might argue that for each of the maximal e-variable there is at least one measure that bring the expectation to one (Grünwald, 2024 calls such property sharpness), sharpness alone is not enough to ensure maximality ($\mathbf{1}_{\mu}$ is a sharp e-variable for $\mathcal{H}_{\mu}$, but it is clearly not maximal).
Two papers (Clerico, 2024; Larsson et al., 2025) have appeared after the first preprint of this work, both extending the results on characterising single-round e-variables for hypotheses defined via linear constraints, a framework that includes the testing for the mean as a special case. Their analyses, however, rely on considerably more abstract and technically advanced proof techniques. Conversely, a key strength of the present paper is to provide a simple and neat argument (illustrated in Figure 1) that works well in the simple setting considered. Interestingly, this same argument is also at the base of the proofs of Theorem 2 and Theorem 3. We also remark that a similar strategy was adopted in the proof of Lemma 2 in Wang (2025). That said, for the multi-round setting, the approach does give rise to non-trivial measurability issues when $\mathcal{X}$ is uncountable. These difficulties already emerge in the two-round case: although one can show that every e-variable must be dominated by a product of the form $(1+\lambda_{1}x_{1})(1+\lambda_{2}(x_{1})x_{2})$, this is constructed in a way that gives no guarantee on the measurability of $\lambda_{2}$. Whether this issue can be resolved remains an open question. Nonetheless, even in the simplified setting we adopt by requiring $\mathcal{X}$ to have finite cardinality [^11] A countable $\mathcal{X}$ would also suffice to solve the measurability issue, though the final step of the proof of Theorem 3 would require more delicate handling., this work is, to the best of the author’s knowledge, the first to fully characterise all e-processes for the natural problem of testing a fixed conditional mean.
@@
To conclude, let us remark that this paper does not aim to delve into the design of effective coin-betting strategies for obtaining tight confidence sequences. For this, we refer the interested reader to the thorough analysis and discussion by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023). Finally, although we have focused on random variables taking values in $[0,1]$ for clarity of exposition, all the results presented readily extend to any bounded closed interval.